a motor pump takes 30 mins to fill tank 2*3*4 m2 which is 10 m height from ground water.find time to fill tank 3*4*5 m2 which is 20 m height from ground water

Grade:12th Pass

1 Answers

Shyama P
11 years ago

Case 1:

Power P=E/t=mgh/t=density*V*g*h/t= (103 * 2*3*4*10*10)/(30*60)=4000/3 Watt

Case 2:

P*t=E  

t (in sec)=(dVgh)/P=(103* 3*4*5*10*20)/(4000/3)=9000sec=150min=2hrs30min